OpenStudy (anonymous):

solve for x & y: 5x+2y=14 6x+3y=24. explain it. thanks. :D

OpenStudy (anonymous):

5x + 2y = 14 6x +3y = 24 -----> it will be ---> y = (24-6x)3 or y = 8 -2x so subtitute it to the first equation 5x + 2(8-2x) = 14 5x + 16 - 4x = 14 ---> add all x and transpose 16 to other side x = 14 - 16 x = -2 ----. substitute it to the 2nd equation 6(-2) + 3y = 24 -12 + 3y =24 3y = 36 y = 12 x = -2 and y = 12